the body starting from rest and travelling with uniform acceleretion covers 15m  during the third second during 5th second how much distance will it cover
 

Grade:9

1 Answers

Vikas TU
6 years ago
Dear Student,
Given that initial velocity is:
u=0, uniform acceleration=a
S=ut+at2/2
=>15=2a+a/2 (Since t=1, for the nth second and u should be the velocity after 2 seconds since we are equating it with the distance travelled in the third second)
=>a=6m/s2

Velocity after 4s = u+at
=0+4a=4a
Distance covered in 5th second=S= ut+at2/2
=>S=4a+a/2
=>S=27m
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)